# Self-Adapting Language Models

### v1

@@ 1 Introduction
As a step towards scalable and efficient adaptation of language models, we propose equipping LLMs with the ability to generate their own training data and finetuning directives for utilizing such data. In particular, we introduce a reinforcement learning algorithm that trains LLMs to generate “self-edits”—natural-language instructions that specify the data and, optionally, the optimization hyperparameters for updating the model’s weights (see Figure 1). We refer to such models as Self-Adapting LLMs (SEAL).
@@ 2 Related Work
Creating synthetic data for training is increasingly common, from large-scale pretraining datasets [15, 16, 17] to task-specific data augmentation [18, 19] and instruction-tuning sets [20, 21]. For incorporation of a smaller-sized corpus, Yang et al. [22] use synthetic data generation via graph-based prompting. SEAL builds on this line of work by using reinforcement learning to train a generative policy that directly maximizes the downstream utility of synthetic data when applied for gradient-based self-updates, rather than relying on static or heuristic generation strategies that are manually tuned and therefore potentially not scalable or optimal.
@@
Several recent works aim to modify or inject factual knowledge into language models via weight updates. Some methods attempt to directly locate specific parameters that correspond to individual facts [23, 24, 25]. Others propose generating additional finetuning data using the information in context [26, 27, 22, 28, 29]. We adopt the latter strategy, following Akyürek et al. [27], who propose generating logical implications of a fact and finetuning on them, and Lampinen et al. [28], who show that implication-based finetuning can even outperform in-context learning. We build on these approaches by training models through RL to generate more optimal finetuning data. Park et al. [29] show that prompting language models to generate question–answer (QA) pairs directly can outperform implication-style prompting. Because the SEAL framework is agnostic to the prompt and format of the self-edit data, it can also be trained to generate QA pairs or other output formats, as explored in §C. With sufficient compute, a cold-start setup, where the model discovers the most effective format without guidance from prompting, may also be viable.
@@
Test-Time Training (TTT) temporarily adapts model weights based on the input the model receives [30, 31, 32, 33]. Akyürek et al. [33] show that combining TTT with ICL enables gradient-updates to outperform standard ICL in the few-shot setting. SEAL can be viewed as incorporating a round of TTT in its inner-loop optimization, leveraging TTT’s efficiency relative to full-scale training to perform multiple updates and reward the generated data that yields the greatest performance gain. Although our method is trained using single-example TTT episodes, we demonstrate in the knowledge incorporation setting that it generalizes to the continued pretraining setting—where placing data directly in context is no longer feasible.
@@
Reinforcement learning has played a central role in improving LLM behavior, originally through RLHF [34]. More recently, RL with verifiable rewards has been applied to boost reasoning performance by optimizing the model directly for task success [35, 36, 37]. SEAL applies RL not to optimize final answers or trace revisions, but to optimize the generation of self-edit data that is then used for weight updates.
@@
SEAL embodies meta-learning principles [38, 39, 40] by learning an adaptation strategy—how to generate effective self-edits—via its outer optimization loop. The goal is to learn how to learn efficiently from task contexts. Meta-learning has similarly been applied in reinforcement learning [41, 42, 43, 44, 45], where models are trained with a meta-objective to rapidly adapt to new tasks. A natural extension of this line of work is self-referential networks, where models modify their own parameters Schmidhuber [46], Irie et al. [47]. In the domain of large language models, recent work has applied meta-learning principles to improve LLM adaptation [48, 49]. Notably, Hu et al. [49] trained a smaller model to output token-specific weights during finetuning on a corpus, addressing a knowledge incorporation task similar to ours. However, SEAL offers greater generality across domains by leveraging the model’s existing generative capabilities to parametrize updates.
@@ 3.1 General Framework
$r(\texttt{SE},\tau,\theta_{t})=\begin{cases}1&\text{If on }\tau\text{, adaptation using }\texttt{SE}\text{ improves }\text{LM}_{\theta_{t}}\text{'s performance}\\ 0&\text{Otherwise}\end{cases}$ (2)
@@ 3.2 Domain Instantiations
Our goal is to efficiently incorporate the information provided in a passage into the model’s weights. A promising recent approach involves using a language model to generate content derived from the passage, followed by finetuning on both the original passage and the generated content [26, 27, 22, 28, 29]. While the form of generated content may vary, we adopt what we consider the canonical format: implications derived from the passage. This approach, introduced in deductive closure training [27], converts a given context $C$ into a set of implications $\texttt{SE}=\{s_{1},s_{2},\dots,s_{n}\}$ by prompting the model to “List several implications derived from the content.” The output may include inferences, logical consequences, or restatements of the original passage. In §C, we also explore alternative prompts such as “rewrite the passage in different ways” or “rewrite in a question-answer format” and show that our method improves performance by similar or greater margins regardless of the base prompt.
@@
The model is prompted with a task’s few-shot demonstrations and generates a self-edit, which in this case is a specification of which tools to invoke and how to configure them, as shown in Figure 3. This self-edit is then applied to adapt the model via LoRA-based finetuning. The adapted model is evaluated on the held-out test input, and the result determines the reward for training the self-edit generation policy.
@@ 4.2 Knowledge Incorporation
Table 2 reports mean no-context SQuAD accuracy under two regimes: single-passage ($n=1$) and continued pretraining (CPT, $n=200$). In the single-passage setting, finetuning directly on the passage yields a negligible gain over the frozen base model (33.5% vs. 32.7%), confirming that using the raw data alone is insufficient. Augmenting with synthetic implications generated by GPT-4.1 boosts accuracy to 46.3%, an improvement of 12.8 percentage points over the passage-only baseline. Using synthetic data produced by the base Qwen-2.5-7B model yields 39.7%, a 6.2-point increase. After reinforcement learning, SEAL further improves accuracy to 47.0%, notably outperforming using synthetic data from GPT-4.1, despite being a much smaller model.
In the CPT setting, the model assimilates information from $n=200$ passages in a single continued pretraining run. It is then evaluated on the union of all 974 corresponding questions. In this setting, we sample $5$ self-edit generations for each passage and take the aggregate synthetic dataset for continued pretraining. As shown in Table 2, we observe a similar ranking of methods as in the single-passage case. SEAL again outperforms all baselines, achieving 43.8% accuracy. While the absolute performance is lower than in the single-passage setting—likely due to increased gradient interference—the relative improvements remain consistent. This suggests that the editing policy discovered by SEAL generalizes beyond its original RL setup of creating synthetic data in a single generation for a single passage.
Method Single Passage (n = 1) Continued Pretraining (n = 200)
Base model 32.7 32.7
Train on Passage 33.5 32.2
Train on Passage + Synthetic 39.7 41.0
Train on Passage + GPT-4.1 Synthetic 46.3 39.4
SEAL 47.0 43.8
Table 2: Knowledge Incorporation Performance across Passage Settings
@@
Figure 5: Example Knowledge Incorporation Self-Edits Across RL Iterations. In this example, we see how RL leads to the generation of more detailed self-edits, which in turn results in better performance. While the progression is clear in this case, the differences across iterations are sometimes more subtle in other examples. We show in §C that prompting for longer self-edits is effective, and that RL training further improves performance by a similar margin.
@@ 5 Limitations
As shown in Figure 6, performance on earlier tasks gradually declines as the number of edits increases, suggesting that SEAL is still susceptible to catastrophic forgetting. Still, it can perform multiple updates without complete collapse, indicating possibility for improvement. Future work could enhance this ability through reward shaping [70, 71, 72] to penalize regressions on earlier tasks, or by integrating continual learning strategies such as null-space constrained edits [73] or representational superposition [74].
@@
Our current instantiations assume that every context is paired with an explicit downstream task: few-shot demonstrations arrive with a held-out query pair, and each passage comes bundled with reference QA. This coupling simplifies reward computation but prevents RL training of SEAL from scaling to unlabeled corpora. A potential solution is to let the model generate not only self-edits but also its own evaluation questions—e.g., draft QA items or synthetic test cases for each passage—while the original content is still in context. These model-written queries could provide the immediate supervision required for reinforcement learning, broadening applicability to general training domains where external question-and-answer sets are unavailable.
@@ 6 Discussion and Conclusion
Villalobos et al. [75] project that frontier LLMs will be trained on all publicly available human-generated text by 2028. We argue that this impending “data wall” will necessitate the adoption of synthetic data augmentation. Once web-scale corpora is exhausted, progress will hinge on a model’s capacity to generate its own high-utility training signal. A natural next step is to meta-train a dedicated SEAL synthetic-data generator model that produces fresh pretraining corpora, allowing future models to scale and achieve greater data efficiency without relying on additional human text.
@@
We would like to thank Shivam Duggal, Idan Shenfeld, Seungwook Han, Jeremy Bernstein, Akarsh Kumar, Linlu Qiu, Juno Kim, Brian Cheung, Moritz Reuss, Ayush Sekhari, Zhang-Wei Hong, Mehul Damani, Leshem Choshen, and Ryan Yang for their valuable discussions and feedback. We acknowledge support from ARO MURI grant number W911NF-23-1-0277. This research was also partly sponsored by the United States Air Force Research Laboratory and the United States Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19- 2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein. This research was also partly supported by the Stevens Fund for MIT UROP research and by the MIT-IBM Watson AI Lab.
@@ A.1 Model and Setup
We use Llama-3.2-1B-Instruct [3] as the base language model. Since this model has no specialized training on ARC, its ability to solve ARC tasks is limited. To enable controlled evaluation, we curated a small set of ARC problems from the training and evaluation splits that are solvable with optimal TTT hyperparameters.
@@ B.6 Standard Error of the Mean in Catastrophic Forgetting Experiment
Appendix C Prompting
Recent works have shown that reinforcement learning baselines and outcomes can be highly sensitive to prompting. We experiment with 4 additional self-edit prompts in the knowledge-incorporation setting. The five prompts—implications, implications-long, implications-very-long, rewrite, and self-qa—are shown below. All results in the main content of the paper used the implications prompt, which we consider to be the most prototypical [27, 28]. However, prior work has found prompts involving rewriting or generating question-answer pairs can be more effective, as discussed in §2.
@@ Appendix C Prompting
No self-edit Implications Implications-long Implications-very-long Rewrite Self-QA
Original 33.5 39.7 49.3 45.0 49.4 37.3
Round 1 43.7 52.4 51.5 55.3 42.8
Round 2 47.0 51.8 52.1 55.6 48.7

### v2

@@ 1 Introduction
As a step towards better language model adaptation, we propose equipping LLMs with the ability to generate their own training data and finetuning directives in response to new inputs. In particular, we introduce a reinforcement learning algorithm that trains LLMs to generate “self-edits”—natural-language instructions that specify the data and, optionally, optimization hyperparameters for updating the model’s weights (see Figure 1). We refer to such models as Self-Adapting LLMs (SEAL).
@@ 2 Related Work
The creation of synthetic data for LLM training is increasingly common, from large-scale pretraining datasets [15, 16, 17, 18, 19] to task-specific data augmentation [20, 21, 22] and instruction-tuning sets [23, 24]. For incorporation of a smaller corpus, Yang et al. [25] use synthetic data generation via graph-based prompting. SEAL builds on this line of work by using reinforcement learning to train a generative policy that directly maximizes the downstream utility of synthetic data when applied for gradient-based self-updates, rather than relying on static or heuristic generation strategies that are manually tuned.
@@
Several recent works aim to modify or inject factual knowledge into language models via weight updates. Some methods attempt to directly locate specific parameters that correspond to individual facts [26, 27, 28]. Others propose generating additional finetuning data using the information in context [29, 30, 25, 31, 32]. We adopt the latter strategy, following Akyürek et al. [30], who propose generating logical implications of a fact and finetuning on them, and Lampinen et al. [31], who show that implication-based finetuning can even outperform in-context learning. We build on these approaches by training models through RL to generate more optimal finetuning data. Park et al. [32] show that prompting language models to generate question–answer (QA) pairs directly can outperform implication-style prompting. Because the SEAL framework is agnostic to the prompt and format of the self-edit data, it can also be trained to generate QA pairs or other output formats, as explored in §B.11.
@@
Test-Time Training (TTT) temporarily adapts model weights based on the input the model receives [33, 34, 35, 36]. Akyürek et al. [36] show that combining TTT with ICL enables gradient-updates to outperform standard ICL in the few-shot setting. SEAL can be viewed as incorporating a round of TTT in its inner-loop optimization, leveraging TTT’s efficiency relative to full-scale training to perform multiple updates and reward the generated data that yields the greatest performance gain. Although our method is trained using single-example TTT episodes, we demonstrate in the knowledge incorporation setting that it generalizes to the continued pretraining (CPT) regime—where placing data directly in context is no longer feasible.
@@
Reinforcement learning has played a central role in improving LLM behavior, originally through RLHF [37, 38]. More recently, RL with verifiable rewards has been applied to boost reasoning performance by optimizing the model directly for task success [39, 40, 41]. SEAL applies RL not to optimize final answers or trace revisions, but to optimize the generation of self-edit data that is then used for weight updates.
@@
SEAL embodies meta-learning principles [42, 43, 44] by learning an adaptation strategy—how to generate effective self-edits—via its outer optimization loop. The goal is to learn how to learn efficiently from task contexts. In reinforcement learning, meta-learning has been used to train agents that learn new tasks quickly [45, 46, 47, 48]. Sun et al. [49] similarly apply RL to learn task-specific weight modulations, offering an alternative to LoRA finetuning that is orthogonal to our approach. A natural extension of meta-learning is self-referential networks, where models modify their own parameters [50, 51]. In the domain of large language models, recent work has applied meta-learning to improve LLM adaptation [52, 53, 54, 55, 49]. Notably, Hu et al. [53] train a smaller model to output token-specific weights during finetuning, addressing a knowledge incorporation task similar to ours, while Chen et al. [54] propose a hypernetwork that generates LoRA adapters conditioned on the input, enabling dynamic and task-specific parameterization. However, SEAL offers greater generality by leveraging the model’s existing generative capabilities to parametrize updates.
@@ 3.1 General Framework
[…]
@@ 3.2 Domain Instantiations
Our goal is to efficiently incorporate the information provided in a passage into the model’s weights. A promising recent approach involves using a language model to generate content derived from the passage, followed by finetuning on both the original passage and the generated content [29, 30, 25, 31, 32]. While the form of generated content may vary, we adopt what we consider the canonical format: implications derived from the passage. This approach, introduced in deductive closure training [30], converts a given context $C$ into a set of implications $\texttt{SE}=\{s_{1},s_{2},\dots,s_{n}\}$ by prompting the model to “List several implications derived from the content.” The output may include inferences, logical consequences, or restatements of the original passage. In §B.11, we also explore alternative prompts such as “rewrite the passage in different ways” or “rewrite in a question-answer format” and show that our method improves performance by similar or greater margins regardless of the base prompt.
@@
The model is prompted with a task’s few-shot demonstrations to generate a self-edit, which in this case is a specification of which tools to invoke and how to configure them, as shown in Figure 3. The self-edit is then applied to adapt the model via LoRA finetuning. The adapted model is evaluated on the held-out test input, and the result determines the reward for the self-edit generation.
@@ 4.2 Knowledge Incorporation
Table 2 reports mean no-context SQuAD accuracy under two regimes: single-passage updating (with LoRA), and small-scale continued pretraining (with full finetuning). We run continued pretraining (CPT) experiments with $n=200$ documents, as well as the full SQuAD validation set of $n=2067$ documents. In the single-passage setting, finetuning directly on the passage yields a negligible gain over the frozen base model (33.5% vs. 32.7%), confirming that using the raw data alone is insufficient. Augmenting with synthetic implications generated by GPT-4.1 boosts accuracy to 46.3%, an improvement of 12.8 percentage points over the passage-only baseline. Using synthetic data produced by the base Qwen-2.5-7B model yields 39.7%, a 6.2-point increase. After reinforcement learning, SEAL further improves accuracy to 47.0%, notably outperforming using synthetic data from GPT-4.1, despite being a much smaller model.
In the CPT setting, the model assimilates information from many passages in a single continued pretraining run. It is then evaluated on the union of all corresponding questions. In this setting, we sample 5 self-edit generations for each passage and take the aggregate synthetic dataset for continued pretraining. As shown in Table 2, we observe a similar ranking of methods as in the single-passage case, but with synthetic data from GPT-4.1 slightly outperforming SEAL. In the $n=200$ setting, SEAL achieves an accuracy of 58.2%, exceeding its single-passage performance. We attribute this gain to the aggregation of multiple self-edit generations. Overall, the strong continued pretraining results of SEAL suggest that the self-editing policy generalizes beyond the original RL setup of creating synthetic data in a single generation for a single passage.
@@
Method Single Passage
(n = 1; LoRA)
Continued Pretraining (n = 200; full-FT) Continued Pretraining
(n = 2067; full-FT)
Base model 32.7 32.7 29.0
Train on Passage 33.5 36.0 31.2
Train on Passage + Synthetic 39.7 50.6 43.4
Train on Passage + GPT-4.1 Synthetic 46.3 59.4 49.2
SEAL 47.0 58.2 46.4
Table 2: Knowledge Incorporation Performance Across Passage Settings.
Figure 5: Example Knowledge Incorporation Self-Edits Across RL Iterations. In this example, we see how RL leads to the generation of more detailed self-edits, which in turn results in better performance. While the progression is clear in this case, the differences across iterations are sometimes more subtle in other examples. We show in §B.11 that prompting for longer self-edits is effective, and that RL training further improves performance by a similar margin.
@@ 5 Limitations
As shown in Figure 6, performance on earlier tasks gradually declines as the number of edits increases, suggesting that SEAL is still susceptible to catastrophic forgetting. Still, it can perform multiple updates without complete collapse, indicating possibility for improvement. Future work could enhance this ability through reward shaping [75, 76, 77] to penalize regressions on earlier tasks, or by integrating continual learning strategies such as null-space constrained edits [78] or representational superposition [79]. In addition, since RL has been shown to forget less than SFT, SEAL’s inner loop could also employ RL instead of SFT [80].
@@
Our current instantiations assume that every context is paired with an explicit downstream task: few-shot demonstrations arrive with a held‑out query pair, and each passage comes bundled with reference QA. This coupling simplifies reward computation but prevents RL training of SEAL from scaling to unlabeled corpora. A potential solution is to let the model generate not only self-edits but also its own evaluation questions—e.g., draft QA items or synthetic test cases for each passage—while the original content is still in context. These model-written queries could provide the immediate supervision required for reinforcement learning, broadening applicability to general training domains where external question-and-answer sets are unavailable.
@@ 6 Discussion and Conclusion
Villalobos et al. [81] project that frontier LLMs will be trained on all publicly available human-generated text by 2028. We argue that this impending “data wall” will necessitate the adoption of synthetic data augmentation. Once web-scale corpora are exhausted, progress will hinge on a model’s capacity to generate its own high-utility training signal. A natural next step is to meta-train a dedicated SEAL synthetic-data generator model that produces fresh pretraining corpora, allowing future models to scale and achieve greater data efficiency without relying on additional human text.
@@
We would like to thank Shivam Duggal, Idan Shenfeld, Seungwook Han, Jeremy Bernstein, Akarsh Kumar, Linlu Qiu, Juno Kim, Brian Cheung, Moritz Reuss, Ayush Sekhari, Zhang-Wei Hong, Mehul Damani, Leshem Choshen, and Ryan Yang for their valuable discussions and feedback. We acknowledge support from ARO MURI grant number W911NF-23-1-0277. This research was also partly sponsored by the United States Air Force Research Laboratory and the United States Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19- 2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein. We acknowledge the MIT Office of Research Computing and Data for providing high performance computing resources that have contributed to the research results reported within this paper. This research was also partly supported by the Stevens Fund for MIT UROP research and by the MIT-IBM Watson AI Lab.
@@ A.1 Model and Setup
For the few-shot learning experiments, we use Llama-3.2-1B-Instruct [3] as the base language model. Since this model has no specialized training on ARC, its ability to solve ARC tasks is limited. To enable controlled evaluation, we curated a small set of ARC problems from the training and evaluation splits that are solvable with optimal TTT hyperparameters.
@@ B.6 Standard Error of the Mean in Catastrophic Forgetting Experiment
B.7 Scaling Model Size
We further experimented with the 3B-parameter Qwen variant, with the same single-passage setup as in Figure 4. The results are given in Table 6.
Table 6: Model Size Scaling Performance (%).
Model Base Model (No Training) Base Model Self-Edit SEAL
Qwen2.5-3B 25.1 31.9 37.0
Qwen2.5-7B 32.7 39.7 47.0
To compare the benefit of SEAL over using self-edits generated by the base model, we compute the ratio of SEAL’s improvement over the base model to the improvement from base model self-edits. This ratio is $1.75\times$ for the 3B model and $2.04\times$ for the 7B model. The relative improvement is greater for the 7B model, which provides some evidence that not only are stronger base models more effective at leveraging synthetic data for self-adaptation, but reinforcement learning may have compounding benefits as model capacity increases. We acknowledge that it is hard to draw conclusions though without actually scaling up further.
B.8 Comparison to Generative Adapter
We additionally compared with Generative Adapter [54], a hypernetwork approach that generates LoRA weights from context, using our evaluation setup. Table 7 reports results for both single-passage (n=1n{=}1) and continued pretraining ($n=200$). We use the Mistral-7B-based model [85] for Generative Adapter, since that was the closest model for comparison. All values are on the same evaluation set, but CPT batches updates over all documents while single-passage trains and evaluates an adapter separately for each document. Generative Adapter achieves strong performance in the n=1n{=}1 case, but underperforms SEAL in the CPT setting. SEAL’s parameterization of weight updates through synthetic data generation allows reuse of generated data for CPT, application to arbitrary base models, and flexibility to learn updates from diverse interaction types beyond LoRA finetuning.
Table 7: SEAL vs. Generative Adapter Performance (%).
Model Base Single-passage (n=1n{=}1) CPT ($n=200$)
SEAL 32.0 47.0 58.2
Generative Adapter 24.4 66.8 28.0
We note that parameterizing weight updates via synthetic data generation rather than directly predicting LoRA weights has several advantages: (1) generated data can be reused for CPT or applied to arbitrary base models, (2) models can leverage reasoning and restructuring as document scale and complexity grow, and (3) the framework is not restricted to LoRA finetuning, allowing for many different update types, including those arising from environment or user interactions. By contrast, it is unclear how hypernetwork-based approaches would scale to such settings, while next-token prediction on generated data naturally exploits a model’s in-context learning capabilities.
B.9 Comparison to Entigraph
We additionally compare SEAL to Entigraph [25] in the Synthetic Continued Pretraining (SCPT) setting on SQuAD. Results for both $200$ and $2067$ passages are shown in Table 8. SEAL uses the same $5$ synthetic data generations per document. For Entigraph, we sample $5$ synthetic data generations involving pairs and $5$ triplets of entities per document. Entigraph with all $10$ synthetic data generations sampling is competitive with SEAL, especially at the larger scale. These results suggest that RL-trained self-edits and structured heuristic methods are both strong approaches for synthetic data generation.
Table 8: Synthetic Continued Pretraining (SCPT) on SQuAD (no passage in context). Best in each column is bolded.
Method Continued Pretraining (n=200) Continued Pretraining (n=2067)
SEAL 58.2 46.4
Entigraph (pairs) 46.2 38.6
Entigraph (pairs+triples) 56.0 48.6
B.10 Proxy Reward
We experiment with replacing the inner loop with a proxy reward based on a human-crafted rubric with 4 categories: length, diversity, quality, and correctness. A GPT-4.1 grader scores each category on a 1-5 scale, and the sum of these scores is used as the RL reward. Table 9 reports final results and RL training times.
Table 9: Full Reward vs. Proxy Reward Performance (%).
Model Base Post-RL Time
SEAL 32.0 47.0 $\approx$6 hr
SEAL w/ Proxy-Reward 32.0 45.6 $\approx$5 min
While further tuning of the rubric or metric design could strengthen the reward signal, the advantage of the full SEAL loop is that no such manual specification is required—the model directly learns which edits improve its own performance. Both approaches appear promising for scaling to larger model sizes and compute budgets: proxy metrics offer dramatically lower cost, and with refinement, they may even surpass the “true” reward of directly optimizing for post-finetuning performance.
B.11 Prompting
Recent works have shown that reinforcement learning baselines and outcomes can be highly sensitive to prompting. We experiment with 6 additional self-edit prompts in the knowledge-incorporation setting. The seven prompts—implications, implications-long, implications-very-long, implications-chain-of-thought, rewrite, self-qa, and no-prompt—are shown below. All results in the main content of the paper used the implications prompt, which we consider to be the most prototypical [30, 31]. However, prior work has found prompts involving rewriting or generating question-answer pairs can be more effective, as discussed in §2.
@@ B.11 Prompting
Method Original Round 1 Round 2 gpt-4.1 synthetic
No self-edit 33.5
Implications 39.7 43.7 47.0 46.3
Implications-long 49.3 52.4 54.4 54.1
Implications-very-long 45.0 51.5 52.1 40.9
Rewrite 49.4 55.3 55.6 54.4
Self-QA 37.3 42.8 48.7 39.2
No-Prompt 13.8 12.7 18.9 28.6
Implications-chain-of-thought 38.7
Chain-of-thought-eval 37.8
@@
Here, “Chain-of-thought-eval” refers to having the model reason before answering the questions (letting the model “pull out” information from its weights), rather than chain-of-thought before generating synthetic data, which is done with the base implications prompt. However, we did not notice a substantial difference in our setting when chain-of-thought was applied, whether before answering and before writing synthetic data.
Letting the model “determine its own” self-edit format, with no-prompt, was not able to achieve the same results as predefined prompting formats in our experiments, achieving only $18.9\%$ after 2 rounds of training.
@@
implications-chain-of-thought Let’s read the following passage, think step by step, and then produce a list of implications derived directly or indirectly from the content. We should first generate a "Thought Process" and then "Implications" Passage: {passage} Thought Process:
no-prompt {passage}